# Benchmarking Quantum Computers: Towards a Standard Performance Evaluation Approach

### v2

@@ III.3.2 Metrics
Gate fidelity: The gate fidelity refers to the accuracy of quantum operations SoKBenchmarkingQuantumComputer . It focuses on how well a quantum gate performs the intended operation compared to the ideal outcome. When a quantum gate is applied, there is a predetermined ideal output, but due to various factors like noise and imperfections in the quantum processor, the actual observed output might deviate from the expected one. High gate fidelity is essential for the reliability and precision of computations. If the gate fidelity is low, it indicates that the quantum gate is introducing errors, reducing the accuracy of quantum algorithms. However, although the concept is clear, its practical application in a quantum processor presents some complex problems. In quantum platforms where the entangling logic gates are two-qubit gates, such as superconducting circuits or photonics, one can apply one of these gates for each pair of connected qubits (two if the gate is asymmetric). Then one can consider metrics such as average fidelity, maximum fidelity, or minimum fidelity. However, it is complicated to compare any of these with other technologies where the gates are multi-qubit, such as trapped ions or cold atoms. In this case, moreover, the fidelity is substantially more costly to calculate.
Decoherence times: The decoherence times refer to the period during which the qubit can maintain its quantum coherence SoKBenchmarkingQuantumComputer . The interaction of qubits with the environment results in the loss of quantum superposition and the decay from the excited state to the ground state, characterized by the times $T_{2}$ and $T_{1}$ respectively. Longer decoherence times are preferable as they might permit deeper circuits to be run, reducing the probability of errors and improving computation accuracy. Unfortunately, coherence times by themselves are not an accurate measure of the quality of the processor. Decoherence occurs due to the coupling of the quantum system to the environment. Strongly protecting a quantum system can result in reduced manipulability and a substantial increase in the time required to operate quantum gates. Consequently, if we double the decoherence times but also double the gate operation times, this has no practical consequence on the processor’s performance, as the number of operations we can carry out will not change. Other thorny question is how to integrate the different coherent times of a qubit and those of different qubits.
Gate speeds: Gate speed is a metric that measures how rapidly a quantum computer can execute a given quantum gate van_der_Schoot_2023 . Considering decoherence times, this metric allows us to determine the number of gates that can be performed before the qubits lose coherence. By increasing gate speeds, we might deploy a greater number of gates within the same time interval, leading to more applied gates during the lifetime of the qubits. Here, the argument explained for coherence times can be inverted, since fast manipulable systems could also be susceptible to decoherence. The problems of this metrics are similar or worse to gate fidelity metric.
@@
Quantum Volume (QV): This is a metric that quantifies the largest square (equal in width and layers) quantum circuit that a quantum processor can successfully run PhysRevA.100.032328 . Each layer consists of random permutation among qubits followed by pseudo-random 2-qubit gates. After running the random quantum circuit, the measured results should be above a certain threshold to be considered successful. Consequently, higher-quality quantum gates and qubits allegedly result in a larger QV. More specifically, ${\rm QV}=2^{D}$, where $D$ is the largest square successfully run. For example, a quantum volume 256 means that a quantum processor can use a subset of 8 qubits in the processor to successfully run a circuit with 8 sequential layers. A higher QV indicates a greater capability of the quantum processor to tackle more complex tasks, demonstrating advancements in the quality of qubits and error mitigation techniques. However, one limitation of QV is that it requires knowing the probabilities of all possible outcomes, which implies a large number of measurements that scale exponentially with the number of qubits. Another limitation is that the QV protocol exclusively considers square circuits. Therefore, QV does not offer a reliable prediction of the performance of circuits that have high depth and low width, or circuits with low depth and high width. Due to the strong limitations of this metric, IBM proposed the alternative Error Per Layered Gate (EPLG) mckay2023benchmarkingquantumprocessorperformance . This metric makes use of (Clifford) randomized benchmarking and mirror circuits techniques to compute the layer fidelity LF, and then $\text{EPLG}=1-\text{LF}^{1/n_{2Q}}$, where $n_{2Q}$ is the number of two-qubit gates, tipically taking the value $N-1$ for a linear chain of qubits. The exponent tries to compensate that the fidelity is a cumbersome measurement of the similarity between two many-body quantum states since it is exponentially suppressed with the size of the processor and, in the thermodynamic limit, two states are either orthogonal or the same. This leads to limitations regarding practicality and reliability, but they will not be analyzed here.
@@
Circuit Layer Operations Per Second (CLOPS): This metric, proposed as a natural extension of Floating Point Operations per Second (FLOPS) in supercomputers, quantifies the number of QV-type layers that the quantum processor can run per second wack2021quality . This metric was proposed by IBM to measure the performance of a processor in terms of speed, complementing QV as a metric, which was focused on measuring quality. Remarkably, CLOPS include also the classical processing and control times in the calculation. Therefore, a larger CLOPS value indicates a higher speed of the quantum processor in executing layers of the same type used in the QV. Therefore, this metric suffers from limitations similar to those of QV regarding scalability, making it neither practical nor repeatable. Additionally, it cannot be adapted to processors making use of multi-qubit gates, such as trapped ions, so it is also not consistent. Finally, as CLOPS include the time corresponding to control, one can improve it by faster classical electronics, without actually improving the processor, so reliability might also be in question. Due to these limitations, IBM proposed an update of this metric called CLOPSh www.ibm.com which updates the definition of layer and now the possible two-qubit gates are no longer arbitrary, but they are only the gates which are native in the processor (i.e. between qubits which are directly connected). The reason is other gates must be decomposed into native ones, jeopardizing the definition of layer as one composed by parallel gates. Although this update is reasonable, it does not fix any of the problems associated to CLOPS.
@@
$\Delta H(p_{\mathrm{exp}})=H(p_{\mathrm{cl}},p)-H(p_{\mathrm{exp}},p),$
where $H(p_{\mathrm{1}},p_{\mathrm{2}})=-\sum^{N}_{i=1}p_{1}(x_{i})\log{p_{2}(x_{i})}$ ) is the cross-entropy between the probabilities distributions $p_{1}$ and $p_{2}$, and $N$ represents the total number of possible outputs. Additionally, the experimental, ideal, and uniform probability distributions are denoted by $p_{\mathrm{exp}}$, $p$, and $p_{\mathrm{cl}}$, respectively. In practice, we will have an experimental implementation of the Haar random unitary $U$ as $A_{\mathrm{exp(U)}}$ with an associated probability distribution $p_{\mathrm{exp}}$ and sample $S_{\mathrm{exp}}=\{x_{1}^{\mathrm{exp}},\dots x_{m}^{\mathrm{exp}}\}$. In this case, the (ideal) experimental cross-entropy difference is defined as
$\alpha\equiv\mathbb{E}_{\mathrm{U}}(\Delta H(p_{\mathrm{exp}})),$
@@
$\alpha\simeq\tilde{\alpha}=H(p_{\mathrm{cl}},p_{\mathrm{U}})-\frac{1}{m}\sum_{j=1}^{m}\log\frac{1}{p_{U}\left(x_{j}^{\mathrm{exp}}\right)},$
with an error of $k/\sqrt{m}$ and $k\simeq 1$. Here, $p_{\mathrm{U}}\left(x_{j}^{\mathrm{exp}}\right)$ is the ideal probability of the $x_{j}^{\mathrm{exp}}$ bitstring computed classically. With $m\sim 10^{3}-10^{6}$ number of samples, a sufficiently accurate approximation can be achieved. The (ideal) experimental cross-entropy difference yields a value ranging from 00 to $1$, which approximates the fidelity of the circuit. When the experimental output probability distribution aligns perfectly with a uniform distribution, it represents the worst-case scenario, obtaining an experimental cross entropy difference equal to zero. On the other hand, when the experimental output distribution perfectly matches the ideal distribution, the experimental cross-entropy difference is one. It is worth emphasizing that experimentally generating random circuits poses a significant challenge. Moreover, obtaining the ideal probability distribution requires classical computational resources that scale exponentially with the number of qubits, making this metric not practical. To discuss whether this metric is repeatable and reliable, it is important to distinguish between the theoretical method used to calculate $\alpha$ and the experimental implementation used to compute $\tilde{\alpha}$. In the theoretical method, the metric satisfies the repeatable and reliable characteristics. In the experimental implementation, there are two options: one can take a fixed finite set of Haar random unitaries, using always this set, or generate a new set of random unitaries every time. The first case satisfies both properties, but the second does not. However, this is more associated with the benchmarking methodology than the metric itself, so we will analyze the metric according to the theoretical method.
@@
$d(P^{\mathrm{ideal}},P^{\mathrm{exp}})=\sqrt{1-BC(P^{\mathrm{ideal}},P^{\mathrm{exp}})},$
@@
$BC(P^{\mathrm{ideal}},P^{\mathrm{exp}})=\sum_{i=1}^{N}\sqrt{p^{\mathrm{ideal}}_{i}p^{\mathrm{exp}}_{i}}.$
The ideal and experimental probabilities to observe the $i$-th outcome are denoted by $p_{i}^{\mathrm{ideal}}$ and $p_{i}^{\mathrm{exp}}$, respectively. The number of possible outcomes is given by $N=2^{n}$, where $n$ is the number of qubits. The Hellinger distance converges to 00 for identical distributions and $1$ in cases, where there is no overlap between distributions. Obtaining the ideal output distribution involves simulating the quantum circuit on a classical computer, which makes this metric not practical. The theoretical method to compute this metric ensures its repeatability and reliability, although its practical application could present challenges.
Heavy Output Generation (HOG): The HOG is a metric that quantifies the success of a pseudo-random quantum circuit executed on a quantum computer 10.5555/3135595.3135617 . If we consider a pseudo-random circuit $U$ taken from a uniform Haar distribution, the ideal and experimental output probability of each bitstring after executing $U$ is denoted by $P^{\mathrm{ideal}}$ and $P^{\mathrm{exp}}$. A bitstring $z$ is classified as heavy in $U$, if its probability $P^{\mathrm{exp}}(z)$ is greater than the median of the distribution $P^{\mathrm{ideal}}$. The probability that a sample drawn from a distribution $P^{\mathrm{exp}}$ is heavy, named as the heavy output generation probability of $P^{\mathrm{exp}}$, is given by
$HOG(P^{\mathrm{exp}},P^{\mathrm{ideal}})=\sum_{i=1}^{N}p^{\mathrm{exp}}_{i}\delta_{i},$
where $N=2^{n}$, $n$ is the number of qubits, and $\delta_{i}=1$ if $i$ is heavy in $U$, and 00 otherwise. If $P^{\mathrm{exp}}$ approximates the uniform distribution, $\mathrm{HOG}(P^{\mathrm{exp}},P^{\mathrm{ideal}})\approx 1/2$; and if $P^{\mathrm{exp}}$ is close to $P^{\mathrm{ideal}}$, $\mathrm{HOG}(P^{\mathrm{exp}},P^{\mathrm{ideal}})\approx(1+\log(2))/2$. The $\mathrm{HOG}$ metric has been employed to develop a test that assesses the performance of a quantum computer as good if it exceeds $2/3$. One limitation of $\mathrm{HOG}$ is that it requires calculating the ideal probability classically, which makes this metric not practical. To determine if this metric is repeatable and reliable, we follow the same criteria used for the cross-entropy difference and XEB metric, which means that it will satisfy both characteristics.
@@
$\ell_{1}(P^{\mathrm{ideal}},P^{\mathrm{exp}})=\sum_{i=1}^{N}|p^{\mathrm{ideal}}_{i}-p^{\mathrm{exp}}_{i}|,$
@@
$\hat{\Delta}\equiv\frac{\hat{R}-N+N(1-e^{-m/N})}{N^{2}/(N+m)-Ne^{-m/N}},$
where $m$ is the number of samples, $N=2^{n}$ with $n$ the number of qubits, and $\hat{R}$ is the observed number of collisions. The expected value of $\hat{\Delta}$ is equal to $1$ when sampling from a typical random quantum state, and 00 when sampling from a uniform distribution. Using this metric, the authors propose a test called collision volume test, which indicates that the computer has good accuracy if $\hat{\Delta}\geq 1/2$. It is important to emphasize that the collision volume metric does not require a classical computer, but it requires around $2^{n/2+5}$ samples to work, which makes it not practical. To determine if this metric is repeatable and reliable, we follow the same criteria used in the cross-entropy difference and XEB metric, which means that it satisfy both quality attributes.

### v3

@@ III.3.2 Metrics
Gate fidelity: The gate fidelity refers to the accuracy of quantum operations. There is no unique definition of fidelity for a quantum operation; however, most definitions are linearly related and share similar properties. Therefore, throughout this text, we refer to fidelity in a general sense. In practical applications, however, it is important that users specify the particular fidelity measure being used. The gate fidelity focuses on how well a quantum gate performs the intended operation compared to the ideal outcome. When a quantum gate is applied, there is a predetermined ideal output, but due to various factors like noise and imperfections in the quantum processor, the actual observed output might deviate from the expected one. High gate fidelity is essential for the reliability and precision of computations. If the gate fidelity is low, it indicates that the quantum gate is introducing errors, reducing the accuracy of quantum algorithms. However, although the concept is clear, its practical application in a quantum processor presents some complex problems. In quantum platforms where entangling gates are natively implemented by two-qubit gates, such as in superconducting circuits or photonic quantum computing, it is common practice to characterize the performance of quantum hardware using metrics associated with two-qubit gates. In such cases, it is natural to evaluate the average, maximum, or minimum fidelity of these gates. However, this logic does not directly apply to other platforms, such as trapped-ion or cold-atom systems, where native multi-qubit gates can be implemented. In these architectures, a single operation can generate collective entanglement among three or more qubits without the need to decompose the operation into a sequence of two-qubit gates. This capability fundamentally change the structure of quantum computation and the efficiency with which certain algorithms can be executed. While comparing individual two-qubit gate fidelities across platforms can be valid when the same operations are implemented, relying solely on such metrics to assess the overall performance of a quantum processor can be misleading. This is particularly true when native gate sets differ significantly, as is the case between platforms supporting multi-qubit gates and those limited to two-qubit gates. In this case, moreover, the fidelity is substantially more costly to calculate. It is also important to note that gate fidelity is not a static property of a quantum hardware. It can vary over time due to coherence times of qubits that are highly sensitive to environmental. Additionally, the fidelity depends on implementation details such as pulse shaping and calibration routines, which can be optimized without manufacture a new quantum hardware.
Readout fidelity: The readout fidelity is a metric that quantifies the accuracy with which a quantum processor measures the final state of qubits. It reflects the probability that the measurement outcome corresponds to the quantum state, typically in the computational basis. High readout fidelity is essential for reliable data collection and accurate measurement of qubit states, especially in NISQ devices, where measurement errors can significantly impact overall performance. Importantly, several techniques can be employed to improve readout fidelity without the need to manufacture a new quantum processor PRXQuantum.5.040326 ; PhysRevApplied.23.054057 ; Seif_2018 .
Decoherence times: The decoherence times refer to the period during which the qubit can maintain its quantum coherence. The interaction of qubits with the environment results in the loss of quantum superposition and the decay from the excited state to the ground state, characterized by the times $T_{2}$ and $T_{1}$ respectively. Longer decoherence times are preferable as they might permit deeper circuits to be run, reducing the probability of errors and improving computation accuracy. Unfortunately, coherence times by themselves are not an accurate measure of the quality of the processor. Decoherence occurs due to the coupling of the quantum system to the environment. Strongly protecting a quantum system can result in reduced manipulability and a substantial increase in the time required to operate quantum gates. Consequently, if we double the decoherence times but also double the gate operation times, this has no practical consequence on the processor’s performance, as the number of operations we can carry out will not change. It is important to note that coherence times are not fixed characteristics of quantum hardware, but time-dependent parameters that may vary due to environmental influences. Additionally, techniques such as dynamical decoupling Viola1998 can be used to improve coherence times by mitigating specific sources of noise applying control pulses to qubits.
Gate speeds: Gate speed is a metric that measures how rapidly a quantum computer can execute a given quantum gate van_der_Schoot_2023 . Considering decoherence times, this metric allows us to determine the number of gates that can be performed before the qubits lose coherence. By increasing gate speeds, a greater number of gates can be executed within the same time interval, leading to more applied gates during the lifetime of the qubits. Here, the argument explained for coherence times can be inverted, since fast manipulable systems could also be susceptible to decoherence. The problems of this metric are similar or worse to gate fidelity metric. It is important to highlight that gate speed is not a fixed metric, as it can be optimized through pulse engineering techniques Werninghaus2021 . Moreover, quantum computing platforms, the strength of the interaction between qubits can be dynamically tuned via external control parameters Yan2018 . Therefore, both pulse engineering techniques and dynamic tuning of qubit interactions provide control over quantum gate speeds.
@@
Quantum Volume (QV): This is a metric that quantifies the largest square (equal in width and layers) quantum circuit that a quantum processor can successfully run PhysRevA.100.032328 . Each layer consists of random permutation among qubits followed by pseudo-random 2-qubit gates. After running the random quantum circuit, the measured results should be above a certain threshold to be considered successful. Consequently, higher-quality quantum gates and qubits allegedly result in a larger QV. More specifically, ${\rm QV}=2^{D}$, where $D$ is the largest square successfully run. For example, a quantum volume 256 means that a quantum processor can use a subset of 8 qubits in the processor to successfully run a circuit with 8 sequential layers. A higher QV indicates a greater capability of the quantum processor to tackle more complex tasks, demonstrating advancements in the quality of qubits and error mitigation techniques. However, one limitation of QV is that it requires knowing the probabilities of all possible outcomes, which implies a large number of measurements that scale exponentially with the number of qubits. Another limitation is that the QV protocol exclusively considers square circuits. Therefore, QV does not offer a reliable prediction of the performance of circuits that have high depth and low width, or circuits with low depth and high width. Due to the strong limitations of this metric, IBM proposed the alternative Error Per Layered Gate (EPLG) mckay2023benchmarkingquantumprocessorperformance . This metric makes use of (Clifford) randomized benchmarking to compute the layer fidelity LF, and then $\text{EPLG}=1-\text{LF}^{1/n_{2Q}}$, where $n_{2Q}$ is the number of two-qubit gates, tipically taking the value N1N-1 for a linear chain of qubits. The exponent is used to convert the aggregate process fidelity of a layer into a “per-gate” error rate by taking the geometric mean. This gives rise to the term Error Per Layered Gate.
@@
Circuit Layer Operations Per Second (CLOPS): This metric, proposed as a natural extension of Floating Point Operations per Second (FLOPS) in supercomputers, quantifies the number of QV-type layers that the quantum processor can run per second wack2021quality . This metric was proposed by IBM to measure the performance of a processor in terms of speed, complementing QV as a metric, which was focused on measuring quality. Remarkably, CLOPS include also the classical processing and control times in the calculation. Therefore, a larger CLOPS value indicates a higher speed of the quantum processor in executing layers of the same type used in the QV. Therefore, this metric suffers from limitations similar to those of QV regarding scalability, making it neither practical nor repeatable. Additionally, the metric does not naturally extend to processors that support native multi-qubit gates, such as trapped-ion hardware, as it assumes a circuit structure composed exclusively of one and two-qubit operations. Consequently, its application to architectures with fundamentally different native gate may lead to inconsistencies in performance comparisons. Finally, as CLOPS include the time corresponding to control, one can improve it by faster classical electronics, without actually improving the processor, so reliability might also be in question. Due to these limitations, IBM proposed an update of this metric called CLOPSh www.ibm.com which updates the definition of layer and now the possible two-qubit gates are no longer arbitrary, but they are only the gates which are native in the processor (i.e. between qubits which are directly connected). The reason is other gates must be decomposed into native ones, jeopardizing the definition of layer as one composed by parallel gates. Although this update is reasonable, it does not fix any of the problems associated to CLOPS.
@@
$\Delta H(p_{\mathrm{exp}})=H(p_{\mathrm{cl}},p)-H(p_{\mathrm{exp}},p),$ (1)
[…]
$\alpha\equiv\mathbb{E}_{\mathrm{U}}(\Delta H(p_{\mathrm{exp}})),$ (2)
@@
$\alpha\simeq\tilde{\alpha}=H(p_{\mathrm{cl}},p_{\mathrm{U}})-\frac{1}{m}\sum_{j=1}^{m}\log\frac{1}{p_{U}\left(x_{j}^{\mathrm{exp}}\right)},$ (3)
with an error of $k/\sqrt{m}$ and $k\simeq 1$. Here, $p_{\mathrm{U}}\left(x_{j}^{\mathrm{exp}}\right)$ is the ideal probability of the $x_{j}^{\mathrm{exp}}$ bitstring computed classically. With $m\sim 10^{3}-10^{6}$ number of samples, a sufficiently accurate approximation can be achieved. The (ideal) experimental cross-entropy difference yields a value ranging from 0 to $1$, which approximates the average gate fidelity of the circuit. When the experimental output probability distribution aligns perfectly with a uniform distribution, it represents the worst-case scenario, obtaining an experimental cross entropy difference equal to zero. On the other hand, when the experimental output distribution perfectly matches the ideal distribution, the experimental cross-entropy difference is one. It is worth emphasizing that experimentally generating random circuits poses a significant challenge. Moreover, obtaining the ideal probability distribution requires classical computational resources that scale exponentially with the number of qubits, making this metric not practical. To discuss whether this metric is repeatable and reliable, it is important to distinguish between the theoretical method used to calculate $\alpha$ and the experimental implementation used to compute $\tilde{\alpha}$. In the theoretical method, the metric satisfies the repeatable and reliable characteristics. In the experimental implementation, there are two options: one can take a fixed finite set of Haar random unitaries, using always this set, or generate a new set of random unitaries every time. The first case satisfies both properties, but the second does not. However, this is more associated with the benchmarking methodology than the metric itself, so we will analyze the metric according to the theoretical method.
@@
$d(P^{\mathrm{ideal}},P^{\mathrm{exp}})=\sqrt{1-BC(P^{\mathrm{ideal}},P^{\mathrm{exp}})},$ (4)
@@
$BC(P^{\mathrm{ideal}},P^{\mathrm{exp}})=\sum_{i=1}^{N}\sqrt{p^{\mathrm{ideal}}_{i}p^{\mathrm{exp}}_{i}}.$ (5)
The ideal and experimental probabilities to observe the $i$-th outcome are denoted by $p_{i}^{\mathrm{ideal}}$ and $p_{i}^{\mathrm{exp}}$, respectively. The number of possible outcomes is given by $N=2^{n}$, where $n$ is the number of qubits. The Hellinger distance converges to 0 for identical distributions and $1$ in cases, where there is no overlap between distributions. Obtaining the ideal output distribution involves simulating the quantum circuit on a classical computer, which makes this metric not practical. The theoretical method to compute this metric ensures its repeatability and reliability, although its practical application could present challenges. Furthermore, it is important to note that the Hellinger distance, like other classical metrics such $\ell_{1}$-Norm Distance, depends on the measurement basis and therefore may fail to capture features such as phase errors.
Heavy Output Generation (HOG):
The (HOG) metric is a measure between two probability distributions, and in the context of quantum computing, it has been used to quantify the success of a pseudo-random quantum circuit executed on a quantum computer 10.5555/3135595.3135617 . If we consider a pseudo-random circuit $U$ taken from a uniform Haar distribution, the ideal and experimental output probability of each bitstring after executing $U$ is denoted by $P^{\mathrm{ideal}}$ and $P^{\mathrm{exp}}$. A bitstring $z$ is classified as heavy in $U$, if its probability $P^{\mathrm{exp}}(z)$ is greater than the median of the distribution $P^{\mathrm{ideal}}$. The probability that a sample drawn from a distribution $P^{\mathrm{exp}}$ is heavy, named as the heavy output generation probability of $P^{\mathrm{exp}}$, is given by
$HOG(P^{\mathrm{exp}},P^{\mathrm{ideal}})=\sum_{i=1}^{N}p^{\mathrm{exp}}_{i}\delta_{i},$ (6)
where $N=2^{n}$, $n$ is the number of qubits, and $\delta_{i}=1$ if $i$ is heavy in $U$, and 0 otherwise. If $P^{\mathrm{exp}}$ approximates the uniform distribution, $\mathrm{HOG}(P^{\mathrm{exp}},P^{\mathrm{ideal}})\approx 1/2$; and if $P^{\mathrm{exp}}$ is close to $P^{\mathrm{ideal}}$, $\mathrm{HOG}(P^{\mathrm{exp}},P^{\mathrm{ideal}})\approx(1+\log(2))/2$. The $\mathrm{HOG}$ metric has been employed to develop a test that assesses the performance of a quantum computer as good if it exceeds $2/3$. One limitation of $\mathrm{HOG}$ is that it requires calculating the ideal probability classically, which makes this metric not practical. To determine if this metric is repeatable and reliable, we follow the same criteria used for the cross-entropy difference and XEB metric, which means that it will satisfy both characteristics.
@@
$\ell_{1}(P^{\mathrm{ideal}},P^{\mathrm{exp}})=\sum_{i=1}^{N}|p^{\mathrm{ideal}}_{i}-p^{\mathrm{exp}}_{i}|,$ (7)
@@
$\hat{\Delta}\equiv\frac{\hat{R}-N+N(1-e^{-m/N})}{N^{2}/(N+m)-Ne^{-m/N}},$ (8)
where $m$ is the number of samples, $N=2^{n}$ with $n$ the number of qubits, and $\hat{R}$ is the observed number of collisions. The expected value of $\hat{\Delta}$ is equal to $1$ when sampling from a typical random quantum state, and 0 when sampling from a uniform distribution. Using this metric, the authors propose a test called collision volume test, which indicates that the computer has good accuracy if $\hat{\Delta}\geq 1/2$. It is important to emphasize that the collision volume metric does not require a classical computer, but it requires around $2^{n/2+5}$ samples to work, which makes it not practical. To determine if this metric is repeatable and reliable, we follow the same criteria used in the cross-entropy difference and XEB metric, which means that it satisfy both quality attributes.